Simplify usinf identites:
186*186*186+14*14*14/186*186-186*14+14*14
1
Expert's answer
2012-06-12T08:48:15-0400
(186*186*186+14*14*14)/(186*186-186*14+14*14)
Let's recall that a3+b3=(a+b)(a2-ab+b2) . In the numerator we have sum of two cubes, in the denominator the second braces from the right part of mentioned formula. Therefore
(186*186*186+14*14*14)/(186*186-186*14+14*14)=(186+14)(186*186-186*14+14*14)/(186*186-186*14+14*14)= =186+14=200